# Saner: Annotation-free Societal Attribute Neutralizer for Debiasing CLIP

### v1

@@ 4.4 Regularization losses
Applying the debiasing loss alone significantly changes original CLIP features, thereby losing semantics [4, 43]. To maintain the alignment of resulting image-text features, we utilize reconstruction loss [43] and contrastive loss [37]. Reconstruction loss $\mathcal{L}_{\text{recon}}$ is the mean squared error between $f_{\text{t}}(t)$ and $h(t)$. Contrastive loss $\mathcal{L}_{\text{cont}}$ aims to minimize the negative log-likelihood of input image-caption pairs, $f_{\text{v}}(v)$ and $f_{\text{t}}(t)$, in comparison to negative ones.
@@ 5 Experiments: Text-to-Image Retrieval
Following previous studies [4, 43, 11, 9], we evaluate SANER on the text-to-image retrieval task regarding gender, age, and racial biases. Further analysis, such as the ablation study of the loss components, is in the supplementary materials.
@@ 5.1 Experimental settings
Evaluation metric. We employ the MaxSkew metric [19], utilized in the previous studies [4, 43, 11, 9], to quantify the societal bias in CLIP in the text-to-image retrieval task. MaxSkew measures the disparity between the attribute distribution of $|\mathcal{A}|$ protected attributes in the top-$k$ retrieved images. Let $\eta_{ak}(q)$ denote the ratio of images labeled with attribute $a$ in the top-$k$ retrieved images. For attribute neutral query $q$, $\eta_{ak}(q)$ should be $1/|\mathcal{A}|$ if the model is unbiased. MaxSkew@$k$ is defined as:
@@
Evaluation setting. For the attribute-neutral queries, we use template-based queries such as “a photo of a $c$ person”, where $c$ is an attribute-neutral concept. Prior work [4] has defined a list of person-related adjectives, such as clever and attractive, as attribute-neutral concepts. We extend the list to encompass occupations (e.g., doctor, nurse), and activities (e.g., cooking and cleaning) for a more comprehensive evaluation. MaxSkew@$k$ is computed per concept. [^10] The complete lists of the concepts are in the supplementary material.
We also evaluate zero-shot image classification accuracy on ImageNet-1K [42] to ensure that debiasing does not spoil the original CLIP’s performance.
Evaluation datasets. We utilize two datasets, FairFace [24] and PATA [43], which consist of images alongside protected attribute annotations (e.g., female and male for gender) associated with the person in each image. FairFace consists of $10,954$ cropped face-centric images, while PATA contains $4,934$ natural images with a single person. Most debiasing approaches only report the performance on the FairFace dataset, but we additionally employ PATA to evaluate the debias performance on more diverse images.
Methods for comparison. We compare SANER against existing methods, i.e., prompt tuning-based debiasing [4] and projection-based debiasing [9], whose code is publicly available. Unfortunately, we could not reproduce the other methods[43, 11] since sufficient reproduction details are unavailable.
Implementation details. Following the previous works [4, 9], we employ CLIP [37] with ViT-B/16 backbone [14] as a target model in our experiments. We train the debiasing layer (a multilayer perception with two linear layers with ReLU activation [34]) appended to the model using $170,624$ image-caption pairs, which is a subset of the COCO training set [29] with person-related words/phrases (e.g., person and boy). We empirically set $\alpha$, $\beta$, and $\gamma$ to $1.0$, $0.1$, and $0.0001$ (Eq. 11), respectively, and train for $5$ epochs. Further details are provided in the supplementary materials.
@@
Original [37] $22.9$ $33.7$ $19.5$ $12.1$ $18.7$ $10.7$
Prompt [4] $12.3$ $29.9$ $20.0$ $6.7$ $16.5$ $10.2$
Projection [9] $15.4$ $37.4$ $15.0$ $6.4$ $13.6$ $5.4$
@@
Original[37] $111.1$ $121.1$ $113.0$ $40.4$ $44.4$ $39.7$
Projection [9] $107.6$ $112.8$ $\mathbf{100.0}$ $37.6$ $45.6$ $45.5$
@@
Original[37] $62.2$ $57.4$ $68.3$ $33.4$ $28.6$ $31.5$
Projection [9] $56.9$ $75.3$ $67.0$ $\mathbf{19.9}$ $43.8$ $26.3$
@@
5.2 Gender bias analysis
Table 1 summarizes MaxSkew@1000 on FairFace and PATA for gender bias, showing that SANER mitigates gender bias the most among all methods. In contrast to the existing methods, this tendency is consistent across 1) the datasets with different image domains (i.e., face-centric and natural images) and 2) concept types (i.e., adjective, occupation, and activity). For instance, while the prompt tuning-based method [4] fails to mitigate bias for activity concepts on FairFace (i.e., bias is amplified from $19.5$ to $20.0$), SANER significantly mitigates bias from $19.5$ to $7.7$. This verifies the better debiasing performance of SANER compared to the existing methods on diverse concepts and image domains, possibly because SANER is trained with diverse text descriptions (i.e., captions in COCO), which are not constrained like pre-defined concepts required in the previous method [4].
Insight: Our debiasing method SANER, which does not necessitate attribute annotations, outperforms previous methods, including those that require annotations, in mitigating gender bias.
5.3 Age and racial biases analysis
Tables 2 and 3 present the results of MaxSkew@1000 for age and racial biases, respectively. We compare SANER with projection-based debiasing [9], as the prompt tuning-based method [4] does not provide age and race debiasing variants. Similarly to the results for gender bias, SANER surpasses the existing method across the datasets and the concept types. For example, SANER successfully mitigates the age bias on PATA. Meanwhile, the projection-based method works only on the adjective concepts and amplifies the bias on the other concepts (e.g., from $39.7$ to $45.5$). These results validate the generalizability of SANER in bias mitigation across the protected attributes.
Insight: SANER also demonstrates superior performance in mitigating age and racial biases compared to the existing method.
5.4 Zero-shot image classification
Table 4: Accuracy on Image-Net-1K.
CLIP Model Acc. (%)
Original [37] $65.4$
Prompt [4] $64.1$
SANER (Ours) $65.2$
To verify whether debiasing harms the zero-shot image classification performance of the original CLIP, we evaluate the prompt tuning-based method [4] and SANER on ImageNet-1K in terms of classification accuracy. Projection-based debiasing [9] does not apply to this evaluation because zero-shot prompts, such as “a photo of a car,” do not necessarily include person-related words. [^11] Projection-based debiasing requires modified input text to include attribute terms. The results are shown in Tab. 4, showing that applying SANER maintains classification performance, whereas the performance of the prompt tuning-based method slightly degrades.
Insight: SANER does not compromise the classification performance of the original CLIP.
6 Experiments: Text-to-Image Generation
We also evaluate SANER on text-to-image generation, for which societal bias is actively investigated [3, 7, 30, 31]. Specifically, we conduct two experiments from different aspects: 1) gender bias regarding occupations using gender-neutral prompts (Sec. 6.1), and 2) retention of attribute information when prompts explicitly disclose gender (Sec. 6.2).
Image generation settings. We use Stable Diffusion (SD) v2.1 [39] as the text-to-image generation model. The CLIP text encoder in SD is replaced with the debiased one for evaluation. Following [9], we use gender-neutral prompts with specifying occupations to analyze gender bias in generated images. These prompts are derived from the template “A photo of a $o$”, where $o$ is replaced with specific occupation terms, such as doctor and teacher. [^12] The list of the occupations is in the supplementary material. On the other hand, we employ gender-specific prompts to evaluate the capability of attribute information retention. Concretely, a gender term, either female or male, is added just before the occupation terms, i.e., “A photo of a {female/male} $o$”, to see if generated images specify the gender. We generate $100$ images for each prompt with SD’s default hyperparameters shown in [39].
Evaluation metrics. For the bias evaluation for the generative task, we use statistical parity (SP) metric that measures the disparity of attribute groups in generated images [8, 49, 9]. Specifically, we annotate binary gender labels (i.e., $\mathcal{A}=\{\texttt{female},\texttt{male}\}$) for the generated images with the assistance of human workers. [^13] Different from previous works that use pre-trained gender classifiers to assign gender labels, we do not use them due to their bias issues [38, 10, 13]. SP is defined as the difference between the empirical distribution $\kappa_{a}$ of gender $a$ and uniform distribution, given by:
6.1 Gender bias analysis
We show visual examples in which SANER successfully mitigates gender bias in Fig. 3. For the gender-neutral prompt, “A photo of a designer,” the original SD and projection-based debiasing (Projection) predominantly generate images depicting a man. In contrast, ours shows a more balanced gender distribution.
Insight: Applying SANER debiased CLIP to Stable Diffusion results in generating images with a more gender-equal distribution.
6.2 Assessment of retention of attribute information
Table 5 also shows the accuracy regarding how much the gender of the person in generated images matches the gender specified in the prompts. For both prompts that describe women and men, using debiased CLIP by projection-based debiasing leads to losing gender information (i.e., accuracies for projection are much lower than those for the original). Conversely, SANER retains gender information (i.e., accuracies for SANER are $1.00$). Figure 4 confirms this, showing that using projection-based debiased CLIP results in generating male images for the prompt, “A photo of a female doctor”.
Insight: SANER retains gender information for input texts that explicitly include gender terms.
Table 5: Gender bias (SP) and retention of gender information (Accuracy) in images generated by Stable Diffusion with original CLIP text encoder (Original) and the model using projection-based debiased CLIP (Projection) and our debiased CLIP (SANER). We report the mean and standard deviation over the set of occupations. Female and male denote prompts with female and male specifications, respectively. Bold represents the best across the models.
Accuracy $\uparrow$
CLIP Model SP $\downarrow$ Female Male
Original [37] $0.51$$\pm 0.22$ $\mathbf{1.00}$$\pm 0.00$ $\mathbf{1.00}$$\pm 0.00$
Projection [9] $0.47$$\pm 0.19$ $0.58$$\pm 0.28$ $0.79$$\pm 0.18$
SANER (Ours) $\mathbf{0.39}$$\pm 0.22$ $\mathbf{1.00}$$\pm 0.00$ $\mathbf{1.00}$$\pm 0.00$
Figure 4: Generated images for the prompt, “A photo of a female doctor,” by the original Stable Diffusion (SD), projection-based debiased CLIP (Projection), and our debiased CLIP (SANER). We randomly sample $10$ images from generated images. Red frame indicates images with incorrect gender (i.e., male).
7 Limitations
Bias metrics. While SANER outperforms the existing methods in mitigating gender, age, and racial biases according to existing bias metrics, such as MaxSkew@$k$ and SP, this does not guarantee that SANER is always superior to them in every aspect or scenario. As suggested in prior work [4, 43, 11, 9], reliance on these metrics provides a standardized measure of societal bias; however, there could be dimensions of bias not adequately captured by them, which implies the necessity of more holistic metrics to assess societal bias.
Further bias mitigation. The experimental results show that SANER noticeably mitigates societal bias in CLIP. Nonetheless, the bias is not completely eliminated (e.g., MaxSkew is not zero). A promising direction for further debiasing could involve debiasing the image encoder, specifically training a debiasing layer to remove attribute information from the visual features for images in which humans do not appear.
Intersectional bias analysis. Although we focus on gender, age, and racial biases separately in our experiments, following the previous debias works [4, 43, 11, 9], SANER can be easily extended to various protected attributes, including their combinations. For instance, considering the intersection of binary gender and age, we generate four sentences with (female, young), (female, old), (male, young), and (male, old) for the debiasing loss, e.g., “A young woman is eating salad” for the input text “A woman is eating salad”. This potential for addressing complex biases is noted in future research.
8 Conclusion
This paper proposed SANER, a simple-yet-effective debiasing method for CLIP, consisting of attribute neutralization and anotation-free debiasing loss. Consequently, SANER can use an arbitrary dataset of image-text pairs for training the debiasing layer, resulting in superior debiasing performance compared to the existing methods for both discriminative and generative tasks (i.e., text-to-image retrieval and text-to-image generation). We also confirmed that SANER retains attribute information for attribute-specific descriptions through the gender-specified prompts for text-to-image generation.
Future work can explore extending SANER to the vision encoder, removing attribute information from visual features in images without humans, and addressing intersectional biases, combining attributes like gender and age. Furthermore, SANER is applicable to any dataset of image-caption pairs, thus, scaling to larger datasets (e.g., Conceptual Captions [44]) for denser supervision is an interesting future direction.
Supplementary Material
This supplementary material includes:
  • Implementation details for SANER (Appendix 0.A).
  • Further analysis (Appendix 0.B).
  • List of person-, gender-, age-, and race-specific terms (Appendix 0.C).
  • List of concepts and occupations (Appendix 0.D).
  • Additional visual examples for image generation experiments (Appendix 0.E).
  • Potential negative impact (Appendix 0.F).
Appendix 0.A Implementation Details for SANER
We train the debiasing layer (a multilayer perception with two linear layers with ReLU activation) using $170,624$ image-caption pairs, which is a subset of the COCO training set with person-related words/phrases defined in Appendix 0.C. The hidden embedding dimensionality for the debiasing layer is set to $128$. We empirically set $\alpha$, $\beta$, and $\gamma$ to $1.0$, $0.1$, and $0.0001$ (Eq. (11) in the main paper), respectively. We set the training epochs, batch size, and learning rate to $5$, $128$, and $5\times 10^{-6}$, respectively.
Racial bias mitigation. For race attribute, we remove race-specific terms [^14] We define race-specific terms. The list is in Appendix 0.C. (e.g., African and Asian) in text descriptions, for instance, “An African woman is eating salad” $\rightarrow$ “A woman is eating salad”.
Appendix 0.B Additional Experiments
0.B.1 Loss ablation
To validate the effectiveness of each regularization loss (i.e., reconstruction loss and contrastive loss), we conduct an ablation study by removing one of the losses or both losses. Table 6 presents the results of gender bias. The results show that using both reconstruction and contrastive losses yields the best results regarding gender bias mitigation and zero-shot classification accuracy on ImangeNet [42]. Furthermore, SANER without regularization losses significantly degrades CLIP’s zero-shot classification ability (i.e., from $65.4$ to $58.8$). These observations confirm the importance of having both reconstruction and contrastive losses for the regularization.
@@ 0.B.1 Loss ablation
0.B.2 Analysis on the data size
The experiments in the main paper (Sections 5 and 6) verify that SANER outperforms existing debiasing methods, showing a better bias mitigation ability in terms of gender and age biases. This superior performance of SANER may be because SANER is trained with diverse text descriptions (i.e., captions in COCO [29]), which are not constrained like pre-defined concepts required in the previous method [4]. In this section, we conduct an experiment to verify this hypothesis. Specifically, we use $n$ percent of the training samples (i.e., $17,624$ image-caption pairs of COCO) to evaluate the impact of the training dataset size. We use the same settings in Appendix 0.A, but use the different training epochs to align the number of iterations. The results are shown in Tab. 7.

### v3

@@ 3.4 Regularization losses
Applying the debiasing loss alone significantly changes original CLIP features, thereby losing semantics (Berg et al., 2022; Seth et al., 2023). To maintain the alignment of resulting image-text features, we utilize reconstruction loss (Seth et al., 2023) and contrastive loss (Radford et al., 2021). Reconstruction loss $\mathcal{L}_{\text{recon}}$ is the mean squared error between $f_{\text{t}}(t)$ and $h(t)$. Contrastive loss $\mathcal{L}_{\text{cont}}$ aims to minimize the negative log-likelihood of input image-caption pairs, $f_{\text{v}}(v)$ and $f_{\text{t}}(t)$, in comparison to negative ones. Note that the original description $t$ is used for regularization losses.
@@ 4 Experiments: Text-to-Image Retrieval
Following previous studies (Berg et al., 2022; Seth et al., 2023; Dehdashtian et al., 2024; Chuang et al., 2023), we evaluate SANER on the text-to-image retrieval task regarding gender, age, and racial biases. Further analysis, such as the ablation study of the loss components, is in the appendix.
@@ 4.1 Experimental settings
Evaluation metric. We employ the MaxSkew metric (Geyik et al., 2019), utilized in the previous studies (Berg et al., 2022; Seth et al., 2023; Dehdashtian et al., 2024; Chuang et al., 2023), to quantify the societal bias in CLIP in the text-to-image retrieval task. MaxSkew measures the disparity between the attribute distribution of $|\mathcal{A}|$ protected attributes in the top-$k$ retrieved images. Let $\eta_{ak}(q)$ denote the ratio of images labeled with attribute $a$ in the top-$k$ retrieved images. For attribute neutral query $q$, $\eta_{ak}(q)$ should be $1/|\mathcal{A}|$ if the model is unbiased. MaxSkew@$k$ is defined as:
@@
Evaluation setting. For the attribute-neutral queries, we use template-based queries such as “a photo of a $c$ person”, where $c$ is an attribute-neutral concept. Prior work (Berg et al., 2022) has defined a list of person-related adjectives, such as clever and attractive, as attribute-neutral concepts. We extend the list to encompass occupations (e.g., doctor, nurse), and activities (e.g., cooking and cleaning) for a more comprehensive evaluation. MaxSkew@$k$ is computed per concept. [^11] The complete lists of the concepts are in the appendix.
We also evaluate zero-shot image classification accuracy on ImageNet-1K (Russakovsky et al., 2015) to ensure that debiasing does not spoil the original CLIP’s performance.
Evaluation datasets. We utilize two datasets, FairFace (Karkkainen & Joo, 2021) and PATA (Seth et al., 2023), which consist of images alongside protected attribute annotations (e.g., female and male for gender) associated with the person in each image. FairFace consists of $10,954$ cropped face-centric images, while PATA contains $4,934$ natural images with a single person. Most debiasing approaches only report the performance on the FairFace dataset, but we additionally employ PATA to evaluate the debias performance on more diverse images.
Methods for comparison. We compare SANER against existing methods, i.e., prompt tuning-based debiasing (Berg et al., 2022) and projection-based debiasing (Chuang et al., 2023), whose code is publicly available. Unfortunately, we could not reproduce the other methods(Seth et al., 2023; Dehdashtian et al., 2024) since sufficient reproduction details are unavailable.
Implementation details. Following the previous works (Berg et al., 2022; Chuang et al., 2023), we employ CLIP (Radford et al., 2021) with ViT-B/16 backbone (Dosovitskiy et al., 2021) as a target model in our experiments. We train the debiasing layer (a multilayer perceptron with two linear layers with ReLU activation (Nair & Hinton, 2010)) appended to the model using $170,624$ image-caption pairs, which is a subset of the COCO training set (Lin et al., 2014) with person-related words/phrases (e.g., person and boy). We empirically set $\alpha$, $\beta$, and $\gamma$ to $1.0$, $0.1$, and $0.0001$ (Eq. 11), respectively, and train for $5$ epochs. Further details are provided in the appendix.
@@
Original (Radford et al., 2021) $22.9$ $33.7$ $19.5$ $12.1$ $18.7$ $10.7$
Prompt (Berg et al., 2022) $12.3$ $29.9$ $20.0$ $6.7$ $16.5$ $10.2$
Projection (Chuang et al., 2023) $15.4$ $37.4$ $15.0$ $6.4$ $13.6$ $5.4$
@@
Table 2: Age bias (left) and racial bias (right), evaluated by MaxSkew@1000 (scaled by $100$), on FairFace. Bold denotes the best across the models. Results on PATA are in the appendix.
FairFace
CLIP Model Adjective Occupation Activity
Original $111.1$ $121.1$ $113.0$
Projection $107.6$ $112.8$ $\mathbf{100.0}$
SANER (Ours) $\mathbf{96.0}$ $\mathbf{112.6}$ $101.9$
FairFace
CLIP Model Adjective Occupation Activity
Original $62.2$ $57.4$ $68.3$
Projection $56.9$ $75.3$ $67.0$
SANER (Ours) $\mathbf{49.3}$ $\mathbf{45.7}$ $\mathbf{46.6}$
4.2 Gender bias analysis
Table 1 presents MaxSkew@1000 on FairFace and PATA for gender bias, showing SANER mitigates bias the most among all methods. In contrast to the existing methods, this tendency is consistent across 1) datasets with different image domains (i.e., face-centric and natural images) and 2) concept types (i.e., adjective, occupation, and activity). For instance, while the prompt tuning-based method (Berg et al., 2022) fails to mitigate bias for activity concepts on FairFace (i.e., bias is amplified from $19.5$ to $20.0$), SANER significantly mitigates bias from $19.5$ to $7.7$. This verifies the better debiasing performance of SANER compared to the existing methods on diverse concepts and image domains, possibly because SANER is trained with diverse text descriptions (i.e., captions in COCO), which are not constrained like pre-defined concepts required in the previous method (Berg et al., 2022).
4.3 Age and racial biases analysis
Tables 2 presents the results of MaxSkew@1000 for age and racial biases. We compare SANER with projection-based debiasing (Chuang et al., 2023), as the prompt tuning-based method (Berg et al., 2022) does not provide age and race debiasing variants. Similarly to the results for gender bias, SANER surpasses the existing method across the datasets and the concept types. For example, SANER successfully mitigates the racial bias on FairFace. Meanwhile, the projection-based method amplifies the bias on the occupation concept (i.e., from $57.4$ to $75.3$). These results validate the generalizability of SANER in bias mitigation across the protected attributes.
4.4 Zero-shot image classification
Table 3: Accuracy on Image-Net-1K.
CLIP Model Acc. (%)
Original $65.4$
Prompt $64.1$
SANER (Ours) $65.2$
To verify whether debiasing harms the zero-shot image classification performance of the original CLIP, we evaluate the prompt tuning-based method (Berg et al., 2022) and SANER on ImageNet-1K in terms of classification accuracy. Projection-based debiasing (Chuang et al., 2023) does not apply to this evaluation because zero-shot prompts, such as “a photo of a car,” do not necessarily include person-related words. [^12] Projection-based debiasing requires modified input text to include attribute terms. The results are shown in Tab. 3, showing that applying SANER maintains classification performance, whereas the performance of the prompt tuning-based method slightly degrades.
Insight: Our method, SANER, which does not require attribute annotations, outperforms previous methods—including those that do require annotations—in mitigating gender, age, and racial biases without compromising classification performance.
5 Experiments: Text-to-Image Generation
We also evaluate SANER on text-to-image generation, for which societal bias is actively investigated (Bansal et al., 2022; Cho et al., 2023; Liu et al., 2024b). Specifically, we conduct two experiments from different aspects: 1) gender bias regarding occupations using gender-neutral prompts (Sec. 5.1), and 2) retention of attribute information when prompts explicitly disclose gender (Sec. 5.2).
Image generation settings. We use Stable Diffusion (SD) v2.1 (Rombach et al., 2022) as the text-to-image generation model. The CLIP text encoder in SD is replaced with the debiased one for evaluation. Following (Chuang et al., 2023), we use gender-neutral prompts with specifying occupations to analyze gender bias in generated images. These prompts are derived from the template “A photo of a $o$”, where $o$ is replaced with specific occupation terms, such as doctor and teacher. [^13] The list of the occupations is in the appendix. On the other hand, we employ gender-specific prompts to evaluate the capability of attribute information retention. Concretely, a gender term, either female or male, is added just before the occupation terms, i.e., “A photo of a {female/male} $o$”, to see if generated images specify the gender. We generate $100$ images for each prompt with SD’s default hyperparameters.
Evaluation metrics. For the bias evaluation for the generative task, we use statistical parity (SP) metric that measures the disparity of attribute groups in generated images (Teo et al., 2023; Chuang et al., 2023). Specifically, we annotate binary gender labels (i.e., $\mathcal{A}=\{\texttt{female},\texttt{male}\}$) for the generated images with the assistance of human workers. [^14] Different from previous works that use pre-trained gender classifiers to assign gender labels, we do not use them due to their bias issues (Ramaswamy et al., 2021; Das et al., 2018; Dinan et al., 2020). SP is defined as the difference between the empirical distribution $\kappa_{a}$ of gender $a$ and uniform distribution, given by:
5.1 Gender bias analysis
Table 4: Gender bias (SP) and gender information retention (Accuracy) in images from Stable Diffusion (Original) and the model using projection-based debiased CLIP (Projection) and our debiased CLIP (SANER). Results are the mean across occupations. Female and male refer to prompts specifying each gender. Bold indicates the best.
Accuracy $\uparrow$
CLIP Model SP $\downarrow$ Female Male
Original $0.51$ $\mathbf{1.00}$ $\mathbf{1.00}$
Projection $0.47$ $0.58$ $0.79$
SANER (Ours) $\mathbf{0.39}$ $\mathbf{1.00}$ $\mathbf{1.00}$
We show visual examples where SANER mitigates gender bias in Fig. 3. For the gender-neutral prompt, “A photo of a designer,” the original SD and projection-based debiasing (Projection) predominantly generate images of a man. In contrast, ours shows a more balanced gender distribution.
5.2 Assessment of retention of attribute information
Table 4 also shows the accuracy of how much the gender of the person in generated images matches the gender specified in the prompts. For both prompts that describe women and men, using debiased CLIP by projection-based debiasing leads to losing gender information (i.e., accuracies for projection are much lower than those for the original). Conversely, SANER retains gender information (i.e., accuracies for SANER are $1.00$). Figure 4 confirms this, showing that using projection-based debiased CLIP results in generating male images for the prompt, “A photo of a female doctor”.
Insight: Applying SANER-debiased CLIP to Stable Diffusion generates images with a more gender-equal distribution while preserving gender information in input texts.
Figure 4: Generated images for the prompt, “A photo of a female doctor,” by the original Stable Diffusion (SD), projection-based debiased CLIP (Projection), and our debiased CLIP (SANER). Red frame indicates images with incorrect gender (i.e., male).
6 Limitations
Further bias mitigation. Our experiments show that SANER noticeably reduces bias in CLIP. Nonetheless, the bias is not completely eliminated (e.g., MaxSkew is not zero). A promising direction for further debiasing could involve debiasing the image encoder, specifically training a debiasing layer to remove attribute information from the visual features for images without human subjects.
Intersectional bias analysis. While our experiments focus on gender, age, and racial biases individually, following prior works (Berg et al., 2022; Seth et al., 2023; Dehdashtian et al., 2024; Chuang et al., 2023), SANER can be easily extended to various protected attributes and their combinations. For instance, considering the intersection of binary gender and age, we generate four sentences with (female, young), (female, old), (male, young), and (male, old) for the debiasing loss, e.g., “A young woman is eating salad” for the input text “A woman is eating salad”. This potential for addressing complex biases is noted in future research.
Use of pre-defined attribute words. While SANER requires general lists of attributes, creating attribute lists is a one-time effort requiring minimal resources, compared to the ongoing cost and complexity of dataset annotation that often needs ethical review and domain expertise.
7 Related Work
As VLMs like CLIP are applied to more tasks, concerns about their social biases have grown (Dehouche, 2021; Wang et al., 2021; Ross et al., 2021; Wolfe et al., 2023; Ruggeri et al., 2023; Srinivasan & Bisk, 2022; Hirota et al., 2024a; b). Hall et al. (Hall et al., 2023) examined gender bias in CLIP, uncovering significant discrepancies in object recognition performance based on the gender depicted in images. such as higher accuracy in recognizing an umbrella with women than with men. These biases risk reinforcing discrimination against marginalized groups (Qiu et al., 2023; Tanjim et al., 2024; Wang et al., 2023). Birhane et al. (Birhane et al., 2021) demonstrated that adopting CLIP-based filtering in the dataset creation can select stereotypical images, like labeling a female astronaut as a ”smiling housewife,” leading to harmful, biased datasets.
8 Conclusion
This paper proposed SANER, a simple-yet-effective debiasing method for CLIP, consisting of attribute neutralization and anotation-free debiasing loss. Consequently, SANER can leverage any image-text dataset to train the debiasing layer, outperforming existing methods in both discriminative and generative tasks (i.e., text-to-image retrieval and text-to-image generation). We also confirmed that SANER retains attribute information for attribute-specific descriptions through the gender-specified prompts for text-to-image generation.
Appendix
This appendix includes:
  • Implementation details for SANER (Appendix A).
  • Further analysis (Appendix B).
  • List of person-, gender-, age-, and race-specific terms (Appendix C).
  • List of concepts and occupations (Appendix D).
  • Additional visual examples for image generation experiments (Appendix E).
  • Potential extension (Appendix F).
  • Potential negative impact (Appendix G).
Appendix A Implementation Details for SANER
We train the debiasing layer (a multilayer perception with two linear layers with ReLU activation) using $170,624$ image-caption pairs, which is a subset of the COCO training set with person-related words/phrases defined in Sec. C. The hidden embedding dimensionality for the debiasing layer is set to $128$. We empirically set $\alpha$, $\beta$, and $\gamma$ to $1.0$, $0.1$, and $0.0001$ (Eq. (11) in the main paper), respectively. We set the training epochs, batch size, and learning rate to $5$, $128$, and $5\times 10^{-6}$, respectively. The training is conducted with a machine equipped with a single NVIDIA A100 GPU 40GB, and it took five hours to train the debiasing layer. Note that the weights of the debiasing layer are updated, and the weights for the rest of the modules are frozen.
Attribute-specific description generation. To avoid computational complexity, we implement a systematic approach using predefined mapping dictionaries between gender terms (e.g., “woman” $\rightarrow$ “man”, “she” $\rightarrow$ “he”). Our algorithm carefully identifies and replaces gender-specific tokens while preserving sentence structure, ensuring no ungrammatical combinations (like “A he/she”) are generated. This ensures efficient and coherent text generation that maintains natural language patterns.
Racial bias mitigation. For race attribute, we remove race-specific terms [^15] We define race-specific terms. The list is in Section C. (e.g., African and Asian) in text descriptions, for instance, “An African woman is eating salad” $\rightarrow$ “A woman is eating salad”.
Appendix B Additional Experiments
B.1 Complete results for age and racial biases
In Table 5 and 6, we show the complete results of Table 2, including results on PATA. The results further verify that SANER demonstrates superior performance in mitigating age and racial biases compared to the existing method.
B.2 Loss ablation
To validate the effectiveness of each regularization loss (i.e., reconstruction loss and contrastive loss), we conduct an ablation study by removing one of the losses or both losses. Table 7 presents the results of gender bias. The results show that using both reconstruction and contrastive losses yields the best results regarding gender bias mitigation and zero-shot classification accuracy on ImangeNet (Russakovsky et al., 2015). Furthermore, SANER without regularization losses significantly degrades CLIP’s zero-shot classification ability (i.e., from $65.4$ to $58.8$). These observations confirm the importance of having both reconstruction and contrastive losses for the regularization.
Negative impact on adding only the contrastive loss. When using only the contrastive loss, MaxSkew scores on FairFace increase compared to no regularization (18.0/19.4/21.7 vs. 15.7/15.0/15.3 without regularization), indicating less effective debiasing. This occurs because the contrastive loss alone focuses on maintaining image-text alignment but does not constrain the debiased features to remain close to the original CLIP features. As a result, the feature modifications may become suboptimal for bias removal. The reconstruction loss plays a crucial role in ensuring the modified features preserve essential semantic information while effectively removing unwanted bias.
@@ B.2 Loss ablation
Original(Radford et al., 2021) $111.1$ $121.1$ $113.0$ $40.4$ $44.4$ $39.7$
Projection (Chuang et al., 2023) $107.6$ $112.8$ $\mathbf{100.0}$ $37.6$ $45.6$ $45.5$
@@
Original(Radford et al., 2021) $62.2$ $57.4$ $68.3$ $33.4$ $28.6$ $31.5$
Projection (Chuang et al., 2023) $56.9$ $75.3$ $67.0$ $\mathbf{19.9}$ $43.8$ $26.3$
@@
B.3 Analysis on the data size
The experiments in the main paper (Sections 5 and 6) verify that SANER outperforms existing debiasing methods, showing a better bias mitigation ability in terms of gender and age biases. This superior performance of SANER may be because SANER is trained with diverse text descriptions (i.e., captions in COCO (Lin et al., 2014)), which are not constrained like pre-defined concepts required in the previous method (Berg et al., 2022). In this section, we conduct an experiment to verify this hypothesis. Specifically, we use $n$ percent of the training samples (i.e., $17,624$ image-caption pairs of COCO) to evaluate the impact of the training dataset size. We use the same settings in Sec. A, but use the different training epochs to align the number of iterations. The results are shown in Table 8.